OpenStudy (anonymous):

solve equation in interval from 0 to 2 pi 7 cos 2t=3

OpenStudy (anonymous):

\[\huge \cos2t=\cos^2 t-\sin^2 t\]

OpenStudy (anonymous):

and I still don't know what to do next

OpenStudy (anonymous):

I rewrote cos2t which means that you should use your logic and use the reqritten expression of cos2t in your equation instead.

OpenStudy (anonymous):

And then \[\huge \cos^2 t -\sin^2 t=2\cos^2 t-1\]

OpenStudy (anonymous):

rewritten*

OpenStudy (anonymous):

I am totally lost

OpenStudy (anonymous):

\[7\cos2t=3\] That is your given equation. Now I rewrite it using what I put up above. \[\large 7(\cos^2 t -\sin^2 t)=3\] Then I rewrite the expression in the brackets again. I'm literally doing this step by step for your own sake and benefit. \[\huge 7(\cos^2 t -(1-\cos^2 t))=3\] I rewrote sin^2 t using my trig identities. \[\cos^2 t + \sin^2 t=1\] Rearrange that trig identity I find myself with this: \[\sin^2 t=1-\cos^2 t\]

OpenStudy (anonymous):

Then \[\huge 7(\cos^2 t-1+\cos^2 t)=3\]

OpenStudy (anonymous):

\[\huge 7(2\cos^2 t -1)=3\]

OpenStudy (anonymous):

Now expand/distribute the LHS.

OpenStudy (anonymous):

14cos^2t-7=3

OpenStudy (anonymous):

14cos^2t=10

OpenStudy (anonymous):

is that right so far

OpenStudy (anonymous):

I don't see where this is getting me. My new equation looks very much like the original one

OpenStudy (anonymous):

That's correct. That's not the original equation. \[\huge 14\cos^2t=10\]

OpenStudy (anonymous):

\[\huge \cos 2t\neq \cos^2 t\]

OpenStudy (anonymous):

Move the 14 to the RHS. by dividing both sides by 14.

OpenStudy (anonymous):

\[\huge \cos^2 t=\frac{10}{14}\]

OpenStudy (anonymous):

Simplify the fraction on the RHS. And then find cost. There will be two solutions since something that's to the power of 2 must have two solutions. One must be negative and the other must be positive.

OpenStudy (anonymous):

5/7

OpenStudy (anonymous):

Yep now find cost.

OpenStudy (anonymous):

by squaring both sides.

OpenStudy (anonymous):

cos of 5/7?

OpenStudy (anonymous):

No. I told you to square root both sides.

OpenStudy (anonymous):

\[\huge \sqrt{\cos^2 t}=\sqrt{\frac{5}{7}}\]

OpenStudy (anonymous):

25/49?

OpenStudy (anonymous):

square root does not equal squaring.

OpenStudy (anonymous):

right

OpenStudy (anonymous):

So you understand that if you square root 1 what would your answer be. This is an example for you to understand that every time you're square rooting something, there should be two solutions. \[\sqrt{1}=?\]

OpenStudy (anonymous):

what would the answer(s) be when you square root 1?

OpenStudy (anonymous):

= and -

OpenStudy (anonymous):

+ and -

OpenStudy (anonymous):

So what would cost equal in this case? \[\cos t=?\]

OpenStudy (anonymous):

\[\huge \cos t=\pm\sqrt{\frac{5}{7}}\]

OpenStudy (anonymous):

Okay now, we find t.

OpenStudy (anonymous):

If there's a +/- sing in front of it, then using logic, there must be 4 solutions for t. One in each quadrant.

OpenStudy (anonymous):

sign*

OpenStudy (anonymous):

So what's t equal to? Get your calculator out.

OpenStudy (anonymous):

I have it

OpenStudy (anonymous):

So did you get t=32*19', 147*41', 212*19' and 327*41'?

OpenStudy (anonymous):

the asterisk is meant to be degrees signs.

OpenStudy (anonymous):

no what did you enter into cal to get those numbers

OpenStudy (anonymous):

I entered: \[\huge \cos^{-1}\sqrt{\frac{5}{7}}=32.3115...\]

OpenStudy (anonymous):

Then I converted that number into degrees and minutes.

OpenStudy (anonymous):

Then I started getting values for the second, third and fourth quadrant.

OpenStudy (anonymous):

so 180-32.311 and 180+32.31 and 360-32.31...

OpenStudy (anonymous):

all those values I got were converted to degrees and minutes

OpenStudy (anonymous):

But if you didn't convert it and just got the values and put the degrees sign up, then you're still correct. But converting it into degrees and minutes is a better answer in maths.

OpenStudy (anonymous):

ok this is really a lot to take in. I need to study this. thanks for doing it step by step. very confusing. do not like this section of algebra

OpenStudy (anonymous):

This is trigonometry for you mate. Work on it. You will get better with time and practice like you said. No worries and good luck.